OpenStudy (anonymous):

how many days are in 4 weeks? in w weeks?

OpenStudy (anonymous):

7*4

OpenStudy (anonymous):

7*4

OpenStudy (anonymous):

7*w= days in w weeks

OpenStudy (anonymous):

1week = 7 days 1 * 4 weeks = 7*4 days 1*w weeks = 7*w days